A barometer kept in a stationary lift reads 76cm .If the elevator starts accelerating up the reading will be

Grade:12th pass

4 Answers

Vikas TU
4 years ago
Hiii 
We know that h = P/row*g
so h is inversly proportional to g 
so, The reading will be less than 76cm.
Hope this helps.
Arun
4 years ago
The pressure of the atmosphere remains the same in both the conditions =ρg*(0.76). But in the upward accelerating elevator the pressure of the mercury column =ρ(g+a)h, since these two are equal so h = (0.76){g/(g+a)}; clearly h
aswanth nayak
4 years ago
dear student,
 
under normal conditions when the elevator is accelerating upwards the reading in the barometer will be less than 76cm (air pressure is same as external) but as acc. to question the pressure is 76 cm means that the air is applying more pressure to lift the liquid to a height of 76cm.
 
Hope it helps you
Regards
aswanth nayak
4 years ago
The pressure of the atmosphere remains the same in both the conditions =ρg*(0.76). But in the upward accelerating elevator the pressure of the mercury column =ρ(g+a)h, since these two are equal so h = (0.76){g/(g+a)}